OpenStudy (anonymous):

a student takes a true false test consisting of 15 questions. assuming that the student guesses at each question, find the probability that the student answers exactly 13 question correctly.

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

the instructions said to use counting rules to determine the probability

OpenStudy (anonymous):

13 ways?

OpenStudy (anonymous):

how can you solve that? i still dont understand

OpenStudy (anonymous):

could you help me find the answer?

OpenStudy (campbell_st):

ok... so here is an alternate method Question choices P (correct answer) 1 2 1/2 2 2 1/2 x 1/2 = 1/4 3 2 1/2 x 1/2 x 1/2 = 1/8 4 2 1/2 x 1/2 x 1/2 x 1/2 = 1/16 . . . . . . the pattern continues..... for 13 correct.... in general its (1/2)^n n is the number of correct answers

OpenStudy (anonymous):

could you please give me the answer? i am not a s good at this as you, even if i try

OpenStudy (campbell_st):

just calculate \[P(13~~correct) =(\frac{1}{2})^{13}\]

OpenStudy (zarkon):

\[{15\choose 13}\left(\frac{1}{2}\right)^{15}\]

OpenStudy (zarkon):

13 right...then they must have 2 wrong